# Manuscript submitted to Aequitas 2024: Workshop on Fairness and Bias in AI | co-located with ECAI 2024, Santiago de Compostela, Spain

### v1

@@ 3.2.2 Non-intersectional Discrimination
The problem with intersectional approaches is that they can lead to an upwards bias in the discrimination score when using meta-metrics [21]. In our case, the number of subgroups grows exponentially with the number of protected attributes. This leads to many subgroups where the number of samples in each subgroup is small, resulting in larger noise in the treatment estimates [21].
Hence, we present a non-intersectional approach, which is more suitable when dealing with a large number of subgroups or when intersectional discrimination is not deemed important. The non-intersectional approach treats each protected attribute independently and aggregates the discrimination scores across all protected attributes. For this, a second aggregate function with $\operatorname{agg}^{(2)}$ is introduced, yielding the following equation:
$\psi_{\text{multi.}}(\mathcal{D})=\underset{Z_{k}\in Z}{\operatorname{agg}^{(2)}}\left\{\underset{i,j\in g_{k},i<j}{\operatorname{agg}^{(1)}}\delta_{Z_{k}}(i,j,E_{1},E_{2})\right\}.$ (12)
The first-level aggregator $\operatorname{agg}^{(1)}$ aggregates disparities within a protected attribute, considering unique pairs of groups $i$ and $j$ in each attribute. The second-level aggregator $\operatorname{agg}^{(2)}$ then combines the results across all protected attributes. By applying both operators, we obtain a discrimination measure that captures disparities between groups across multiple attributes.
@@ 3.2.3 Example
$\psi_{\text{intersect.}}(\mathcal{D})$ $=\underset{i,j\in G_{\text{avail}}}{\max}\hat{\delta}_{Z}(i,j,(Y=1),\text{True})$ (13)
$=\underset{i,j\in G_{\text{avail}}}{\max}\hat{\delta}_{\text{\{Age, Sex\}}}(i,j,(Y=1),\text{True})$ (14)
$=\underset{i\in G_{\text{avail}}}{\max}P(Y=1\mid Z_{1}=i_{1},Z_{2}=i_{2})-\underset{j\in G_{\text{avail}}}{\min}P(Y=1\mid Z_{1}=j_{1},Z_{2}=j_{2})$ (15)
$=|P(Y=1\mid\text{Age}=\text{Old},\text{Sex}=\text{Male})-P(Y=1\mid\text{Age}=\text{Young},\text{Sex}=\text{Male})|=1.$ (16)
We show in the following that the discrimination for the non-intersectional approach is different as it treats each protected attribute independently:
$\psi_{\text{multi.}}(\mathcal{D})$ $=\underset{Z_{k}\in Z}{\max}\left\{\underset{i,j\in g_{k},i<j}{\max}\delta_{Z_{k}}(i,j,(Y=1),\text{True})\right\}$ (17)
$=\max\left\{\delta_{\text{Age}}(\text{Old},\text{Young},(Y=1),\text{True}),\delta_{\text{Sex}}(\text{Male},\text{Female},(Y=1),\text{True})\right\}$ (18)
$=\max\{|0.5-0.5|,|0.5-0.5|\}=\max\{0,0\}=0.$ (19)
The non-intersectional approach yields a discrimination score of 0 because the disparities for both protected attributes are 0. This is quite different from the intersectional approach, which reports a discrimination score of 1. As seen, the choice of the approach can lead to different results.
@@ 4 Experiments
Our experimentation follows a pipeline consisting of data pre-processing, bias mitigation, model training, and evaluation. To mitigate bias in tabular datasets with multiple protected attributes, we used a sampling method, FairDo [9], that constructs fair datasets by selectively sampling data points. The method is very flexible and only requires the user to define the discrimination measure that should be minimized. In our case, we are interested in yielding a fair dataset that minimizes the bias for multiple protected attributes. The experiments revolve around two research questions:
@@ 4.1 Experimental Setup
The tabular datasets used in our experiments are: Adult [14], Bank [15], and Compas [16]. They are known for their use in fairness research and contain multiple protected attributes. We pre-processed the datasets by applying one-hot encoding to categorical variables and label encoding to protected attributes. Table 2 shows important characteristics of the datasets after pre-processing.
@@
Applying the bias mitigation method FairDo [9] to the datasets can be regarded as a pre-processing step, too. This is because the method simply returns a dataset that is fair with respect to the given discrimination measure. FairDo [9] offers a variety of methods to mitigate bias and we chose the undersampling method that removes samples. The fairness objective is stated as [9]:
@@
where $\mathcal{D}$ is the training set of Adult, Bank, or Compas, and $\psi$ is the fairness objective function. We considered $\psi_{\text{intersect.}}$ and $\psi_{\text{multi.}}$ as our objectives. The first measures intersectional discrimination, while the latter measures discrimination for multiple protected attributes where the intersection of groups is not considered. Bias mitigation is only applied to the training set and the testing set remains unchanged. FairDo internally uses genetic algorithms for optimization. We used the same settings and operators as provided in the package and only adjusted the population size (200) and the number of generations (400).
@@
We utilized the scikit-learn library [26] to train various machine learning classifiers, namely Logistic Regression (LR), Support Vector Machines (SVM), Random Forest (RF), and Artificial Neural Networks (ANNs). These classifiers were trained on both the original and fair datasets. Classifiers trained on the original datasets serve as a baseline for comparison. We used the default hyperparameters given by scikit-learn for each classifier.
@@
We evaluated the models’ predictions on fairness and performance using the test set. For fairness, we assessed $\psi_{\text{intersect.}}$ and $\psi_{\text{multi.}}$. For the classifiers’ performances, we report the area under the receiver operating characteristic curve (AUROC) [27], where higher values indicate better performance. Because removing data points can compromise the overall quality of the data, we also report the number of subgroups before and after bias mitigation, as well as the number of samples.
@@
For each dataset, the bias mitigation process was repeated 10 times. The results were averaged over the trials to obtain a more robust evaluation.
@@ 4.2 Results
Disc. Before Disc. After Subgroups Before Subgroups After
Dataset Metrics
Adult $\psi_{\text{multi.}}$ $20\%$ $13\%$ 10 10
$\psi_{\text{intersect.}}$ $31\%$ $16\%$ 10 10
Bank $\psi_{\text{multi.}}$ $24\%$ $5\%$ 48 48
$\psi_{\text{intersect.}}$ $33\%$ $15\%$ 48 46.2
Compas $\psi_{\text{multi.}}$ $30\%$ $5\%$ 34 34
$\psi_{\text{intersect.}}$ $100\%$ $17\%$ 34 28.8
Table 3 shows the average discrimination before and after mitigating bias in the training sets. On all datasets, discrimination was reduced after applying FairDo. It can be observed that calculating the discrimination for multiple protected attributes while considering intersectionality leads to a higher discrimination value. Without considering group intersections, discrimination was reduced by 7%, 19%, and 25% for Adult, Bank, and Compas, respectively. When considering intersectionality, the discrimination was reduced by 15%, 18%, and 83%. Hence, discrimination was reduced by 28% on average across all datasets.
We also report the number of subgroups before and after bias mitigation to assess the impact of the undersampling method on the dataset. The removal of subgroups can only be observed in the intersectional setting. In the Compas dataset 5.2 out of 34 subgroups were removed on average, indicating the largest amount of subgroups removed across all datasets. While the Bank dataset consists of 48 subgroups, only 1.8 subgroups were removed on average. Because the Compas dataset’s initial intersectional discrimination score is 100%, removing multiple subgroups seems inevitable to reduce discrimination.
@@
Figure 2: Results on the test set. The x-axis represents the discrimination values (legend indicates used measure) and the y-axis represents the classifiers’ performances. We compare the pre-processed (fair) data with the original data. The points represent the average and the error bars show the standard deviation of the AUROC and discrimination values over 10 trials.
Figure 2 shows the results of the classifiers’ performances on the test set. The classifiers’ performances are displayed on the y-axis, while the discrimination values are shown on the x-axis. We show AUROC to evaluate the classifiers’ performances, where higher values indicate better performances. We note that the axes are not shared across the subfigures for better visibility.
Classifiers trained on fair datasets did not suffer a significant decline of performance than those trained on original datasets. In all cases, only a slight decrease of 1%-3% in performance can be observed. This indicates that the bias mitigation process does not compromise the dataset’s fidelity and therefore, the classifiers’ performances. Regarding discrimination, a great reduction can be seen. The x-axis scales are much larger than the y-axis scales, suggesting that changes in discrimination are larger than changes in performance. For example, the RF classifier trained on the Bank dataset (Figure 2g) shows a decrease in intersectional discrimination from 38% to 15%, while the performance only decreases by 2%. The results suggest that FairDo can be reliably used to mitigate bias in tabular datasets for various measures that consider multiple protected attributes. Still, we advise users to carefully perform similar analyses when applying the method to their datasets.
@@ 5 Discussion
The results of our experiments show that the presented measures detect discrimination in datasets with multiple protected attributes differently. When using the intersectional discrimination measure, more groups are identified and compared to each other, which leads to a higher reported discrimination score. While subgroups are not ignored by this measure, exponentially more groups are considered, and measuring higher discrimination scores by random chance becomes more likely [20, 21]. In contrast, treating each protected attribute separately prevents this issue but may lead to an underestimation of discrimination. The choice of measure is up to the stakeholders and depends on the context of the dataset. It also depends on the regulations for the application of the AI system. We generally recommend using the intersectional discrimination measure if the interplay of protected attributes is of interest and the number of groups is low. If a higher number of groups is present, the number of individuals in each subgroup should be high enough to prevent high discrimination scores that are due to noise. Otherwise, treating each protected attribute separately is more suitable.
Using the mitigation strategy FairDo [9], the resulting datasets in the experiments have improved statistical properties regarding fairness. Whether intersectionality was considered or not, reducing discrimination in datasets was possible, which led to training fairer machine learning models without compromising their performances substantially. At the current state, the AI Act [1] does not explicitly mention intersectional discrimination nor how to deal with multiple protected attributes generally. While recital (67) states that datasets “should […] have the appropriate statistical properties”, it does not specify what these properties are. Hence, our work serves as an initial guideline for what these properties could be and how to achieve them in practice.
@@ 6 Conclusion
Datasets often come with multiple protected attributes, which makes measuring and mitigating discrimination more challenging. Most existing studies only deal with a single protected attribute, and works that consider multiple protected attributes often focus on intersectionality. In opposition to this, we have proposed a new measure that treats each protected attribute separately. This is more suitable when the number of subgroups is too high and the number of individuals in each subgroup is low. We used both measures in our experiments and used them as objectives for the bias mitigation method FairDo. On average, the disparity between protected groups was reduced by 28% in the bias-mitigated datasets compared to the original datasets. Machine learning models trained on these datasets also showed a reduction in discrimination while maintaining their performance.

### v3

@@ 3.2.2 Non-intersectional Discrimination
The problem with measuring discrimination for intersectional groups is that it has an upward bias when using meta-metrics [22]. This is because the number of subgroups grows exponentially with the number of protected attributes. This leads to many subgroups where the number of samples in each subgroup is possibly small, resulting in larger noise in the treatment estimates [22].
Besides intersectional groups, Yang et al. [11] listed a non-intersectional definition of groups, called independent groups. Building on the definition of independent groups, we propose an appropriate approach to measure discrimination for this type of groups. It is more suitable when dealing with a large number of subgroups or when intersectional discrimination is not deemed important. Our non-intersectional approach treats each protected attribute independently and aggregates the discrimination scores across all protected attributes. For this, a second aggregate function with $\operatorname{agg}^{(2)}$ is introduced, yielding the following equation:
$\psi_{\text{indep}}(\mathcal{D})=\underset{Z_{k}\in\mathcal{Z}}{\operatorname{agg}^{(2)}}\left\{\underset{i,j\in g_{k},i<j}{\operatorname{agg}^{(1)}}\delta_{Z_{k}}(i,j,E_{1},E_{2})\right\}.$ (11)
The first-level aggregator $\operatorname{agg}^{(1)}$ aggregates disparities within a protected attribute, considering unique pairs of groups $i$ and $j$. The second-level aggregator $\operatorname{agg}^{(2)}$ then combines the results across all protected attributes. By applying both operators, we obtain a discrimination measure that captures disparities between groups across multiple attributes.
@@ 3.2.3 Example
$\psi_{\text{intersect}}(\mathcal{D})$ $=\underset{i,j\in G_{\text{avail}}}{\max}\hat{\delta}_{\mathcal{Z}}(i,j,(Y=1),\Omega)$ (12)
$=\underset{i,j\in G_{\text{avail}}}{\max}\hat{\delta}_{\text{\{Age, Sex\}}}(i,j,(Y=1),\Omega)$
$=\underset{i\in G_{\text{avail}}}{\max}P(Y=1\mid Z_{1}=i_{1},Z_{2}=i_{2})-\underset{j\in G_{\text{avail}}}{\min}P(Y=1\mid Z_{1}=j_{1},Z_{2}=j_{2})$
$=|P(Y=1\mid\text{Age}=\text{Old},\text{Sex}=\text{Male})-P(Y=1\mid\text{Age}=\text{Young},\text{Sex}=\text{Male})|=1,$
while the discrimination score for the non-intersectional approach is given by:
$\psi_{\text{indep}}(\mathcal{D})$ $=\underset{Z_{k}\in Z}{\max}\left\{\underset{i,j\in g_{k},i<j}{\max}\delta_{Z_{k}}(i,j,(Y=1),\Omega)\right\}$ (13)
$=\max\left\{\delta_{\text{Age}}(\text{Old},\text{Young},(Y=1),\Omega),\delta_{\text{Sex}}(\text{Male},\text{Female},(Y=1),\Omega)\right\}$
$=\max\{|0.5-0.5|,|0.5-0.5|\}=\max\{0,0\}=0.$
The non-intersectional approach yields a discrimination score of 0 because the disparities for both protected attributes are 0. This is quite different from the intersectional approach, which reports a discrimination score of 1. As seen, the results can differ depending on the approach.
@@ 4 Experiments
Our experimentation follows a pipeline consisting of data pre-processing, bias mitigation, model training, and evaluation. To mitigate bias in tabular datasets with multiple protected attributes, we used the sampling method, FairDo [9], that constructs fair datasets by selectively sampling data points. The method is very flexible and only requires the user to define the discrimination measure that should be minimized. In our case, we are interested in a dataset that has minimal bias across multiple protected attributes. The experiments revolve around the following research questions:
@@ 4.1 Experimental Setup
The tabular datasets employed in our experiments include the Adult [15], Bank [16], and COMPAS [17] datasets. They are known for their use in fairness research and contain multiple protected attributes. We pre-processed the datasets by applying one-hot encoding to categorical variables and label encoding to protected attributes. Table 2 shows important characteristics of the datasets after pre-processing.
@@
Applying the bias mitigation method FairDo [9] to the datasets can be regarded as a pre-processing step, too. This is because the method simply returns a dataset that is fair with respect to the given discrimination measure. FairDo [9] offers a variety of options to mitigate bias, and we chose the undersampling method that removes samples. In this option, the optimization objective is stated as [9]:
@@
where $\mathcal{D}$ is the training set of Adult, Bank, or COMPAS, and $\psi$ is the fairness objective function. We experimented with both $\psi_{\text{intersect}}$ and $\psi_{\text{indep}}$ as objectives functions. Bias mitigation is only applied to the training set and the testing set remains unchanged. FairDo internally uses genetic algorithms to select a subset of the training set that minimizes the objective function. We used the same settings and operators as provided in the package and only adjusted the population size (200) and the number of generations (400).
@@
We utilized the scikit-learn library [27] to train various machine learning classifiers, namely Logistic Regression (LR), Support Vector Machine (SVM), Random Forest (RF), and Artificial Neural Network (ANN). These classifiers were trained on both the original and fair datasets. Classifiers trained on the original datasets serve as a baseline for comparison. We used the default hyperparameters given by scikit-learn package for each classifier.
@@
We evaluated the models’ predictions on fairness and performance using the test set. For fairness, we assessed $\psi_{\text{intersect}}$ and $\psi_{\text{indep}}$. For the classifiers’ performances, we report the area under the receiver operating characteristic curve (AUROC) [28], where higher values indicate better performances. Because removing data points can compromise the overall quality of the data, we also report the number of subgroups before and after bias mitigation to check for representativeness.
@@
For each dataset and discrimination measure combination, the bias mitigation process was repeated 10 times. The results were averaged over the trials to obtain a more robust evaluation.
@@ 4.2 Results
Dataset Metric Disc. Before Disc. After Subgroups Before Subgroups After
Adult $\psi_{\text{indep}}$ $20\%$ $13\%$ 10 10
$\psi_{\text{intersect}}$ $31\%$ $16\%$ 10 10
Bank $\psi_{\text{indep}}$ $24\%$ $5\%$ 48 48
$\psi_{\text{intersect}}$ $33\%$ $15\%$ 48 46.2
COMPAS $\psi_{\text{indep}}$ $30\%$ $5\%$ 34 34
$\psi_{\text{intersect}}$ $100\%$ $17\%$ 34 28.8
Table 3 shows the average discrimination before and after mitigating bias in the training sets. On all datasets, discrimination was reduced after applying FairDo. Without considering group intersections, discrimination was reduced by 7%, 19%, and 25% for Adult, Bank, and COMPAS, respectively. When considering intersectionality, the discrimination was reduced by 15%, 18%, and 83%. Hence, discrimination was reduced by 28% on average across all datasets, thus answering RQ1 positively. When comparing the discrimination scores, it can be observed that the intersectional discrimination scores are generally higher. This is because in the intersectional setting, more subgroups are considered, which potentially leads to larger differences between them [21].
We also report the number of subgroups before and after bias mitigation to assess the impact of the undersampling method on the dataset. The removal of subgroups can only be observed in the intersectional setting. In the COMPAS dataset 5.2 out of 34 subgroups were removed on average, indicating the largest amount of subgroups removed across all datasets. While the Bank dataset consists of 48 subgroups, only 1.8 subgroups were removed on average. Because the COMPAS dataset’s initial intersectional discrimination score is 100%, removing more subgroups seems inevitable to reduce bias.
@@
Figure 2: Results on the test set. The x-axis represents the discrimination values (legend indicates used measure) and the y-axis represents the classifiers’ performances. We compare the pre-processed (fair) data with the original data. The points/stars represent averages, and the error bars display the standard deviations of the AUROC and discrimination values over 10 trials.
Figure 2 shows the results of the classifiers’ performances on the test set. The classifiers’ performances are displayed on the y-axis, while the discrimination values are shown on the x-axis. We note that the axes do not share the same scale across the subfigures for analytical purposes.
Classifiers trained on fair datasets did not suffer a significant decline in performance compared to those trained on original datasets. In all cases, only a slight decrease of 1%-3% in performance can be noted. This indicates that the bias mitigation process does not compromise the dataset’s fidelity and, therefore, the classifiers’ performances. Regarding discrimination, a significant reduction is evident. The x-axis scales are much larger than the y-axis scales, suggesting that changes in discrimination are larger than changes in performance. For example, the RF classifier trained on the Bank dataset (Figure 2g) shows a decrease in intersectional discrimination from 38% to 15%, while the performance only decreases by 2%. Similar results can be observed for the other classifiers and datasets as well, successfully addressing RQ2. The results suggest that FairDo can be reliably used to mitigate bias in tabular datasets for various measures that consider multiple protected attributes. Still, we advise users to carefully perform similar analyses when applying the method to their datasets.
@@ 5 Discussion
The results of our experiments show that the presented measures detect discrimination in datasets with multiple protected attributes differently. When using the intersectional discrimination measure, more groups are identified and compared to each other. While subgroups are not ignored by this measure, measuring higher discrimination scores by random chance becomes more likely [21, 22]. In contrast, treating each protected attribute separately prevents this issue but may lead to overlooking discrimination. The choice of measure is up to the stakeholders and depends on the context of the dataset and the regulations that apply to the AI system. We generally recommend using the intersectional discrimination measure if the number of individuals in each subgroup is large enough to draw statistically significant conclusions. Otherwise, treating each protected attribute separately is more suitable.
By using the mitigation strategy FairDo [9], the resulting datasets in the experiments have improved statistical properties regarding fairness. Whether intersectionality was considered or not, reducing discrimination in datasets was possible. At the current state, the AI Act [1] does not explicitly mention intersectional discrimination nor how to deal with multiple protected attributes generally. While recital (67) states that datasets “should […] have the appropriate statistical properties”, it does not specify what these properties are. Hence, our work serves as an initial guideline for what these properties could be and how to achieve them in practice.
@@ 6 Conclusion
Datasets often come with multiple protected attributes, which makes measuring and mitigating discrimination more challenging. Most existing studies only deal with a single protected attribute, and works that consider multiple protected attributes often focus on intersectionality. In opposition to this, we proposed a new non-intersectional measure that treats each protected attribute separately. This is more suitable when the number of subgroups is too large or the number of individuals in each subgroup is small. We used both intersectional and non-intersectional measures as objectives and applied the FairDo framework to mitigate discrimination in multiple datasets. The experiments show that discrimination was reduced in all datasets and on average by 28%. Machine learning models trained on the bias-mitigated datasets also improved their fairness while maintaining performance compared to models trained on the original datasets.